# Provably Safe Generative Sampling with Constricting Barrier Functions

### v1

@@ 7 Conclusion
We introduced a framework for enforcing hard constraints on flow-based generative models by framing guided sampling as an optimal control problem. Our approach leverages constricting Control Barrier Functions (CBFs) to define a safety tube that is relaxed at the initial noise distribution and progressively constricts to the target safe set. The resulting minimum-norm QP synthesizes a feedback control input that provably maintains samples within the safety tube while minimizing the instantaneous contribution to the distributional shift from the learned model, as quantified by the KL divergence (Theorem 4.2). We validated the modularity and effectiveness of this framework across physically-consistent trajectory generation, constrained image synthesis, and smooth action generation for robotic manipulation, using off-the-shelf pre-trained models. As generative models are increasingly deployed in safety-critical systems, this framework provides a deterministic safety layer that complements the expressiveness of generative sampling models.

### v2

@@ 7 Conclusion
We introduced a framework for enforcing hard constraints on flow-based generative models by framing guided sampling as a problem of control synthesis. Our approach leverages constricting Control Barrier Functions (CBFs) to define a safety tube that is relaxed at the initial noise distribution and progressively constricts to the target safe set. The resulting minimum-norm QP synthesizes a feedback control input that provably maintains samples within the safety tube (Theorem 4.1) while minimizing the instantaneous contribution to the distributional shift from the learned model, as quantified by the KL divergence (Theorem 4.2). We validated the modularity and effectiveness of this framework across physically-consistent trajectory generation, constrained image synthesis, and smooth action generation for robotic manipulation, using off-the-shelf pre-trained models. Our framework maintains semantic fidelity of the model while ensuring constraint adherence. As generative models are increasingly deployed in safety-critical systems, this framework provides a deterministic safety layer that complements the expressiveness of generative sampling models.